Lisaneedham:

i need help The value of a collector’s item is expected to increase exponentially each year. The item is purchased for $500 and its value increases at a rate of 5% per year. Find the value of the item after 4 years. $578.81 $607.75 $1687.50 $2531.25

supie:

I think what you have to do is \[500 × 1.21550625 = ?\]

Lisaneedham:

thanks

supie:

np

jimthompson5910:

Agreed. You would do \(y = P*(1+r)^x\) \(y = 500*(1+0.05)^4\) \(y = 500*(1.05)^4\) \(y = 500*1.21550625\) \(y = ??\) Where P = 500 is the initial value, r = 0.05 is the growth rate, and x = 4 is the number of years.